A 6.20mL sample of 0.10mol/hydrochloride acid is used to standardized sodium hydroxide. What the concentration of the sodium hydroxide?
1
Expert's answer
2020-06-12T07:30:17-0400

The reaction between hydrochloric acid and sodium hydroxide is as following:


"NaOH + HCl \rightarrow NaCl + H_{2}O;"


Thus, we can calculate the amount in moles of HCl in a given solution:


"n(HCl)=C(HCl)*V(HCl)=0.10mol\/L*0.00620L=6.2*10^{-4}mol;"


The reactants are interacting in a ratio 1 to 1. Thus, the amount of sodium hydroxide is equal to 6.2*10-4 moles.